OMG Inc. has 5 million shares of common stock outstanding, 4 million shares of preferred stock outstanding, and 6,000 bonds. Suppose the common shares sell for $16 per share, the preferred shares sell for $15 per share, and the bonds sell for 109 percent of par.

What weight should you use for preferred stock in the computation of OMG’s WACC? (Round your answer to 2 decimal places.)
Answer #1

Total value of:

Common stock=(5*16)=$80 million

Preferred stock=(4*15)=$60 million

Bonds=6000*(1000*109%)=$6.54 million

Total value=(80+60+6.54)=146.54 million

Hence weight of preferred stock=(60/146.54)

=40.94%(Approx).